# Neural Multigrid Memory For Computational Fluid Dynamics

### v1

@@ Abstract
Turbulent flow simulation plays a crucial role in various applications, including aircraft and ship design, industrial process optimization, and weather prediction. In this paper, we propose an advanced data-driven method for simulating turbulent flow, representing a significant improvement over existing approaches.
Our methodology combines the strengths of Video Prediction Transformer (VPTR) (Ye & Bilodeau, 2022) and Multigrid Architecture (MgConv, MgResnet) (Ke et al., 2017). VPTR excels in capturing complex spatiotemporal dependencies and handling large input data, making it a promising choice for turbulent flow prediction. Meanwhile, Multigrid Architecture utilizes multiple grids with different resolutions to capture the multiscale nature of turbulent flows, resulting in more accurate and efficient simulations.
Through our experiments, we demonstrate the effectiveness of our proposed approach, named MGxTransformer, in accurately predicting velocity, temperature, and turbulence intensity for incompressible turbulent flows across various geometries and flow conditions. Our results exhibit superior accuracy compared to other baselines, while maintaining computational efficiency.
@@ 4.2 Rayleigh–Bénard convection (RBC)
This is a model for turbulent convection, with a horizontal layer of fluid heated from below so that the lower surface is at a higher temperature than the upper surface, governed by Navier-Stokes Equations.
The dataset comes from two dimensional turbulent flow simulated using the Lattice Boltzmann Method. We use only the velocity vector fields, where the spatial resolution of each frame is $1,792\times 256$ (see Figure [2]). Each image has two channels, one is the turbulent flow velocity along $x$ direction and the other one is the velocity along $y$ direction. The physics parameters relevant to this numerical simulation are: Prandtl number = $0.71$, Rayleigh number = $2.5\texttimes 10^{8}$ and the maximum Mach number = $0.1$. We use $1,000$ images for our experiments. The task is to predict the spatiotemporal velocity fields up to $50$ steps ahead given $16$ initial frames.
@@
The 2D Random Flow dataset utilized in this experiment is derived from the PDEBench benchmark (Takamoto et al., 2022), a comprehensive resource for machine learning experiments in physics (Takamoto et al., 2022).
For this particular experiment, we focused on the random flow dataset within the 2D Computational Fluid Dynamics (CFD) sections. The dataset was generated through simulations of the compressible Navier-Stokes equation, considering a shear viscosity of $\eta=0.01$ and a bulk viscosity of $\zeta=0.01$. These parameters enable the modeling of fluid behavior in a computationally efficient manner.
The dataset comprises a total of 10,000 samples, each of which contains $21$ frames. Each frame consists of two channels representing the $x$-velocity and $y$-velocity components of the fluid flow. The resolution of each frame is set to $128\times 128$, providing a sufficiently detailed representation of the flow characteristics.
To ready the dataset for testing, we resized the samples to dimensions of $(21,2,64,64)$, where the values indicate the number of timesteps, channels, image height, and image width, respectively. Subsequently, we divided these downscaled samples into three sets: 3,000 for training, 1,000 for validation, and 1,000 for testing. For our model evaluations, we performed tests on 5,000 of these test samples across all models.
@@ 4.4 Result and Observation
Table 1 presents the evaluation metrics, namely Mean Squared Error (MSE) and Divergence, for a single prediction step. In terms of the one-step prediction performance, MGxTransformers demonstrates strong potential by outperforming most of the baselines. However, when employing autoregressive methods, our model falls short compared to the TF-Net and FNO2d baselines. Furthermore, regarding divergence loss, our models exhibit a slight performance deficit. This can be attributed to the absence of normal physical laws within the model’s structure. The lack of physical reflection within the architecture is the underlying cause of this limitation.
Figure [3] visually represents the evolution of the evaluation metrics over a span of $100$ prediction steps, employing an autoregressive method. Notably, our model exhibits a diminishing performance over time, primarily attributed to its inability to preserve essential physical properties and image details within a single prediction step. Nonetheless, these observations underscore the potential efficacy of incorporating physical regularizers as a promising avenue for enhancing our model in future iterations.
@@ 5 Conclusion
We introduce a fully data-driven approach for learning the behavior of turbulent flow. Our methodology incorporates a Transformer layer that effectively captures the spatio-temporal interactions inherent in turbulent flow, akin to a variant of Conv-LSTM. Additionally, the inclusion of MgNet layer facilitates the simulation of the multiscale structure characteristics of fluid dynamics. The end-to-end training pipeline aims to achieve the optimal combination of these components.
To comprehensively evaluate the performance of our proposed model, MGxTransformers, we conduct extensive comparisons against various baselines, employing different metrics to assess their respective strengths and limitations. A significant contribution of this research lies in the novel concept of modeling the multiscale structure of turbulent flow through purely deep learning methods.
Moving forward, our future work entails the incorporation of physical regularizers such as divergence and turbulence kinetic energy. Moreover, we aim to explore methods that directly integrate the modeling of multiscale structures within the Transformer layers. These endeavors collectively aim to enhance the accuracy and fidelity of deep learning models in the context of turbulent flow analysis.

### v2

@@ Abstract
Turbulent flow simulation plays a crucial role in various applications, including aircraft and ship design, industrial process optimization, and weather prediction. In this paper, we propose an advanced data-driven method for simulating turbulent flow, representing a significant improvement over existing approaches. Our methodology combines the strengths of Video Prediction Transformer (VPTR) (Ye & Bilodeau, 2022) and Multigrid Architecture (MgConv, MgResnet) (Ke et al., 2017). VPTR excels in capturing complex spatiotemporal dependencies and handling large input data, making it a promising choice for turbulent flow prediction. Meanwhile, Multigrid Architecture utilizes multiple grids with different resolutions to capture the multiscale nature of turbulent flows, resulting in more accurate and efficient simulations. Through our experiments, we demonstrate the effectiveness of our proposed approach, named MGxTransformer, in accurately predicting velocity, temperature, and turbulence intensity for incompressible turbulent flows across various geometries and flow conditions. Our results exhibit superior accuracy compared to other baselines, while maintaining computational efficiency. Our implementation in PyTorch is available publicly at https://github.com/Combi2k2/MG-Turbulent-Flow.
@@ 4 Experiments
Our source code is available at https://github.com/Combi2k2/MG-Turbulent-Flow.
@@ 4.2 Rayleigh–Bénard convection (RBC)
This is a model for turbulent convection, with a horizontal layer of fluid heated from below so that the lower surface is at a higher temperature than the upper surface, governed by Navier-Stokes Equations. The dataset comes from two dimensional turbulent flow simulated using the Lattice Boltzmann Method. We use only the velocity vector fields, where the spatial resolution of each frame is $1,792\times 256$ (see Figure [2]). Each image has two channels, one is the turbulent flow velocity along $x$ direction and the other one is the velocity along $y$ direction. The physics parameters relevant to this numerical simulation are: Prandtl number = $0.71$, Rayleigh number = $2.5\texttimes 10^{8}$ and the maximum Mach number = $0.1$. We use $1,000$ images for our experiments. The task is to predict the spatiotemporal velocity fields up to $50$ steps ahead given $16$ initial frames.
@@
The 2D Random Flow dataset utilized in this experiment is derived from the PDEBench benchmark (Takamoto et al., 2022), a comprehensive resource for machine learning experiments in physics (Takamoto et al., 2022). For this particular experiment, we focused on the random flow dataset within the 2D Computational Fluid Dynamics (CFD) sections. The dataset was generated through simulations of the compressible Navier-Stokes equation, considering a shear viscosity of $\eta=0.01$ and a bulk viscosity of $\zeta=0.01$. These parameters enable the modeling of fluid behavior in a computationally efficient manner. The dataset comprises a total of 10,000 samples, each of which contains $21$ frames. Each frame consists of two channels representing the $x$-velocity and $y$-velocity components of the fluid flow. The resolution of each frame is set to $128\times 128$, providing a sufficiently detailed representation of the flow characteristics. To make the dataset ready for testing, we resized the samples to dimensions of $(21,2,64,64)$, where the values indicate the number of timesteps, channels, image height, and image width, respectively. Subsequently, we divided these downscaled samples into three sets: 3,000 for training, 1,000 for validation, and 1,000 for testing. For our model evaluations, we performed tests on 5,000 of these test samples across all models.
@@ 4.4 Result and Observation
Table 1 presents the evaluation metrics, namely Mean Squared Error (MSE) and Divergence, for a single prediction step. In terms of the one-step prediction performance, MGxTransformers demonstrates strong potential by outperforming most of the baselines. However, when employing autoregressive methods, our model falls short compared to the TF-Net and FNO2d baselines. Furthermore, regarding divergence loss, our models exhibit a slight performance deficit. This can be attributed to the absence of normal physical laws within the model’s structure. The lack of physical reflection within the architecture is the underlying cause of this limitation. Figure 3 visually represents the evolution of the evaluation metrics over a span of $100$ prediction steps, employing an autoregressive method. Notably, our model exhibits a diminishing performance over time, primarily attributed to its inability to preserve essential physical properties and image details within a single prediction step. Nonetheless, these observations underscore the potential efficacy of incorporating physical regularizers as a promising avenue for enhancing our model in future iterations.
@@ 5 Conclusion
We introduce a fully data-driven approach for learning the behavior of turbulent flow. Our methodology incorporates a Transformer layer that effectively captures the spatio-temporal interactions inherent in turbulent flow, akin to a variant of Conv-LSTM. Additionally, the inclusion of MgNet layer facilitates the simulation of the multiscale structure characteristics of fluid dynamics. The end-to-end training pipeline aims to achieve the optimal combination of these components. To comprehensively evaluate the performance of our proposed model, MGxTransformers, we conduct extensive comparisons against various baselines, employing different metrics to assess their respective strengths and limitations. A significant contribution of this research lies in the novel concept of modeling the multiscale structure of turbulent flow through purely deep learning methods. Moving forward, our future work entails the incorporation of physical regularizers such as divergence and turbulence kinetic energy. Moreover, we aim to explore methods that directly integrate the modeling of multiscale structures within the Transformer layers. These endeavors collectively aim to enhance the accuracy and fidelity of deep learning models in the context of turbulent flow analysis.